# On multi-graded Proj schemes

### v3

@@ 5.5 Derived categories
on derived categories. This functor sends objects in $D^{?}_{\mathrm{neg}}{\mathsf{Mod}}^{M}(A)$ to complexes all of whose cohomology objects are trivial, i.e. to the zero object, so that our functor factors through a triangulated functor
$D^{?}{\mathsf{Mod}}^{M}(A)/D^{?}_{\mathrm{neg}}{\mathsf{Mod}}^{M}(A)\to D^{?}\mathsf{QCoh}({\operatorname{Proj}}^{M}(A)).$
To check that this functor is an equivalence, we note that by [Mi91, Theorem 3.2] there exists a canonical equivalence of triangulated categories
@@ 5.6 On a lemma by Arkhipov–Bezrukavnikov
As in Remark 5.14, under the present assumptions, the functor $Q\mapsto\widetilde{Q}$ identifies with the functor
@@
[…]
$\phi:{\operatorname{Hom}}_{K^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M}(A)/K_{\mathrm{neg}}^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M}(A)}({\mathcal{F}},{\mathcal{F}}^{\prime})\xrightarrow{\sim}{\operatorname{Hom}}_{D^{\mathrm{b}}\mathsf{Coh}^{T}({\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger}))}(j^{*}{\mathcal{F}},j^{*}{\mathcal{F}}^{\prime}).$
@@
$\mathcal{F}\otimes_{A}\mathcal{C}[-1]\to{\mathcal{F}}$
with the same property. We claim that the composition
@@
vanishes. This will imply that the composition of the first two morphisms factors through a morphism
$\mathcal{F}\otimes_{A}\mathcal{C}[-1]\to{\mathcal{F}}^{\prime}$
(in $D^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg}}^{M}(A)$, or equivalently in $K^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M}(A)$); then the diagram
@@
    Consider the setting of Remark 5.4, with $H$ a linearly reductive algebraic group over $𝕜\Bbbk$. Then one can consider the full subcategory ${\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M,H}(A)$ of the category ${\mathsf{Mod}}_{\mathrm{fg}}^{M,H}(A)$ of objects which sums of objects of the form $V\otimes A(\alpha)$ where $V$ is a finite-dimensional $H$-module and $\alpha\in M$. Let $K_{\mathrm{neg}}^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M,H}(A)$ be the full triangulated subcategory of $K^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M,H}(A)$ whose objects are the complexes all of whose cohomology objects are supported on $V(A_{\dagger})$. Then the same proof as for Proposition 5.29 shows that the natural functor

### v4

@@ 5.5 Derived categories
on derived categories. This functor sends objects in $D^{?}_{\mathrm{neg}}{\mathsf{Mod}}^{M}(A)$ to complexes all of whose cohomology objects are trivial, i.e. to the zero object, so that our functor factors through a triangulated functor $D^{?}{\mathsf{Mod}}^{M}(A)/D^{?}_{\mathrm{neg}}{\mathsf{Mod}}^{M}(A)\to D^{?}\mathsf{QCoh}({\operatorname{Proj}}^{M}(A))$. To check that this functor is an equivalence, we note that by [Mi91, Theorem 3.2] there exists a canonical equivalence of triangulated categories
@@ 5.6 On a lemma by Arkhipov–Bezrukavnikov
[…]
@@
where $j:{\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\to{\operatorname{Spec}}(A)$ is the embedding. What we have to show is therefore that for any bounded complexes ${\mathcal{F}},{\mathcal{F}}^{\prime}$ of objects in ${\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M}(A)$ the morphism
$\phi:{\operatorname{Hom}}_{K^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M}(A)/K_{\mathrm{neg}}^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M}(A)}({\mathcal{F}},{\mathcal{F}}^{\prime})\to{\operatorname{Hom}}_{D^{\mathrm{b}}\mathsf{Coh}^{T}({\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger}))}(j^{*}{\mathcal{F}},j^{*}{\mathcal{F}}^{\prime}).$
@@
vanishes. This will imply that the composition of the first two morphisms factors through a morphism $\mathcal{F}\otimes_{A}\mathcal{C}[-1]\to{\mathcal{F}}^{\prime}$ (in $D^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg}}^{M}(A)$, or equivalently in $K^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M}(A)$); then the diagram
@@
    Consider the setting of Remark 5.4, with $H$ a linearly reductive algebraic group over $𝕜\Bbbk$. Then one can consider the full subcategory ${\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M,H}(A)$ of the category ${\mathsf{Mod}}_{\mathrm{fg}}^{M,H}(A)$ of objects which are sums of objects of the form $V\otimes A(\alpha)$ where $V$ is a finite-dimensional $H$-module and $\alpha\in M$. Let $K_{\mathrm{neg}}^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M,H}(A)$ be the full triangulated subcategory of $K^{\mathrm{b}}{\mathsf{Mod}}_{\mathrm{fg},\mathrm{fr}}^{M,H}(A)$ whose objects are the complexes all of whose cohomology objects are supported on $V(A_{\dagger})$. Then the same proof as for Proposition 5.29 shows that the natural functor